OpenStudy (anonymous):

transform this equation into slope intercept form: 4y - 8 = - 6x

OpenStudy (anonymous):

transform this equation into slope intercept form: 4y - 8 = - 6x

OpenStudy (anonymous):

the slope intercept form is y =mx +c

OpenStudy (anonymous):

so try to arrange the given equation and keep the term containing y on the left side of equal sign and all other term on the right side of the = sign

OpenStudy (anonymous):

and then try to make the coefficient of y as 1 by dividing both sides by the coefficient of y

OpenStudy (anonymous):

4y - 8 = - 6x 4y=-6x +8 y = (-6x +8)/4 y= -6x/4 +8/4 y=-3x/2 +2 slope is =-3/2 and y-intercept = +2

OpenStudy (anonymous):

OH MY GOD THANK YOU SO MUCH! You are my hero:-)